OpenStudy (anonymous):

Write an equation of the line that passes through (8, 3) and (4, 3). A. y = 3 B. y = 4x + 3 C. y = 4 D. y = 3x

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (tkhunny):

Do you have a 2-point form?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

first use the slope formula m = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

there is no slope

OpenStudy (anonymous):

@jim_thompson5910 actually the y-coordinates are same in both cases hence the equation will be y = 3

jimthompson5910 (jim_thompson5910):

well you shouldn't give the answer away first of all, and second, the slope is 0 since y = 3 is the same as y = 0x + 3 m = 0 (slope) b = 3 (y-intercept)

jimthompson5910 (jim_thompson5910):

to find that the slope is 0, you would plug in the given points into the slope formula m = (y2-y1)/(x2-x1)